# A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6

### v2

@@ 4 Discussion
In section 3, we have reported the properties of polarised dust in SPT034652, which indicates that kpc-scale ordered magnetic fields are already present at the Cosmic Dawn. Here, we focus on several key questions that can be addressed from these observations: (1) How different are the magnetic fields between the distant and local galaxies? (2) What are the main drivers in ordering galactic magnetic fields? (3) What is the formation timescale of magnetic fields? (4) How could the early large-scale magnetic fields affect galaxy evolution?
@@ 4.1 Polarisation fraction and morphology of magnetic fields
Owing to the wavelength coverage and the spatial resolution provided by ALMA, we can observe similar rest-frame FIR and reach comparable spatial resolutions as the dust polarisation observations in the nearby Universe, especially compared with what High-resolution Airborne Wideband Camera Plus (HAWC+) has done in the SALSA Legacy Program (Lopez-Rodriguez et al. 2022b). In Fig. 4, we compare the measured dust polarisation fraction in SPT0346-52 with the measurements from the SALSA Legacy Program. The rest-frame wavelength probed in SPT034652 is at 131 m, averaged over a physical resolution about 700 pc, comparable to the spatial resolution of SALSA Legacy Program at 154 m (250 pc-1.5 kpc, Lopez-Rodriguez et al. 2022b). A typical mean resolved polarisation fraction of per cent has been reported among all the galaxies in the SALSA survey, dropping to a lower fraction of per cent if considering only starburst galaxies (SFR). We also include the dust polarisation measurements of 9io9 (Geach et al. 2023). The similar polarisation fraction between SPT034652 and the local starbursts is surprising, even though there are two orders of magnitude differences in SFR and more than 12 Gyr difference in cosmic time.
In SPT034652, the polarised dust emission is more patchy. We only observe polarised dust emission in a part of the galaxy and, interestingly, it does not overlap with the peak of total dust intensity (see Fig. 2). In Fig. 5, we compare the spatial distribution of polarisation with the kinematics from [C ii]. The region with the polarisation detection is also the region with the lowest velocity gradient and velocity dispersion. In the channel maps of [C ii] emission, the polarisation is mostly associated with a single velocity component near . This result is different from the configuration of polarised dust emission in 9io9, where the polarised dust is largely overlapped with the total dust emission and the inferred fields line are aligned with cold molecular gas disk (Geach et al. 2023). Such a difference suggests a possible different ordering process (see more discussion in section 4.2).
@@ 4.2 Physical origin of the ordered magnetic field
Various processes could lead to the formation of large-scale ordered magnetic fields, including the dynamo in galactic disks (Beck et al. 1994; Beck 2007), the stretching and compressing of existing fields by galaxy mergers, and galactic inflows/outflows (Chyży & Beck 2004; Drzazga et al. 2011; Basu et al. 2017; Jones et al. 2019b; Lopez-Rodriguez et al. 2022a; Lopez-Rodriguez 2023; Ledos et al. 2024). In this section, we will discuss several possible mechanisms to explain the observed magnetic fields in SPT034652.
4.2.1 Disk or merger driven?
@@
High-resolution spectral line observations of SPT034652 have resolved a large velocity gradient ( 600 km s-1) of the cold gas in the ISM (Spilker et al. 2015; Litke et al. 2019). Due to the compact size of SPT034652 in the source plane, it is still unclear whether it’s a rotating disk or an ongoing major merger. Litke et al. (2019) showed that a lensing reconstruction of the high-resolution [C ii] emission revealed two distinct components, both spatially and kinematically. However, based on the observation of water absorption, Jones et al. (2019a) concluded that a compact disk with a strong molecular outflow gives a better explanation. In this subsection, we discuss both scenarios in generating the ordered magnetic fields.
@@
The formation of disks is the most discussed channel to develop galactic ordered magnetic fields (Beck et al. 1994; Pakmor et al. 2017, 2024). The presence of a galactic disk can provide continuous differential rotation to support the dynamo (Beck et al. 2019). Initially, dynamo was not favoured in the early Universe, as most of the early galaxies were thought to be irregular and turbulent. This picture may need to be revised, however, given the apparent prevalence of disks in the early Universe (Förster Schreiber et al. 2018; Rizzo et al. 2023; Kuhn et al. 2023, Lee et al. (in prep.)). If galaxies can develop disks efficiently during the first Gyr of the cosmic time, it is reasonable to expect large-scale ordered magnetic fields in the disks of early galaxies (Arshakian et al. 2009; Geach et al. 2023).
@@
In SPT034652, the rich observations allow a quantitative diagnosis of dynamo associated with a galactic disk. Following the simplified model by Arshakian et al. (2009), we can define the dynamo number:
@@
where the is the scale height of the disk, is the angular velocity of the disk; and is the maximum velocity and velocity dispersion of the disk; is the maximum radius of the disk. Within a galactic disk, the dynamo is effective only when the dynamo number is larger than the critical value , which is close to 7 in Milky Way-like disk galaxies. The from our lens model is about 1.5 kpc (see also Fig. 3). Assuming a typical disk scale height of 300 pc (Geach et al. 2023), the is in the order of . Considering the complex velocity field (merger+outflows), it is not straightforward to estimate the without some biased assumptions. In the most extreme case, assuming the velocity gradient is solely produced by a galactic disk, the maximum rotational velocity will be km s-1. We adopted the median velocity dispersion of 140 km s-1 from the region with the lowest velocity gradient, which is also the region with polarisation detection (see also Fig. 5). Then, the is about 2.5, which predicts , well below the (Ruzmaikin et al. 1988; Arshakian et al. 2009, see also the discussion in Geach et al. (2023)). Therefore, even if there is a disk in SPT034652, it will not be very effective in ordering the galactic magnetic fields. However, if the dynamic cold disk () could exist at such early cosmic time (e.g. Rizzo et al. 2020; Neeleman et al. 2020; Lelli et al. 2021; Rowland et al. 2024), the dynamo number could be large enough to allow efficient dynamo.
@@
Galaxy mergers can compress and stretch magnetic fields, creating large-scale ordered magnetic fields. This process has been well demonstrated in the local major merger prototype – NGC4038/39 (the Antennae) (Chyży & Beck 2004; Basu et al. 2017). In the cores of the two merging galaxies of the Antennae system, the field lines are mostly dominated by turbulent fields. However, in the tidal tail, the magnetic fields are regular up to kpc, which is likely generated by stretching the initial disk fields in the pre-merger galaxies (Basu et al. 2017). A 9 kpc ordered magnetic field is also reported by means of magnetically aligned dust grains, which also supports a compressed and stretched magnetic field due to the merger activity (Lopez-Rodriguez et al. 2022a). The presence of ordered magnetic fields has also been confirmed in a sample of interacting galaxies observed with synchrotron polarisation (at 4.86 and 1.4 GHz) (Drzazga et al. 2011), which all support tidal interactions can effectively magnetise the mergers’ surroundings and produce large-scale ordered field structures up to tens of kpc in the ISM of the merged galaxies. Besides, the developed kpc-scale ordered magnetic fields have a strength several times larger than the magnetic fields in disks from non-interacting galaxies (Drzazga et al. 2011; Beck & Wielebinski 2013). Similar intense field amplification and creation of large-scale magnetic fields with galaxy mergers have also been noticed in magneto-hydrodynamic simulations, which is crucial for angular momentum transportation and subsequent morphological evolution of the merged galaxies (Whittingham et al. 2021, 2023).
@@
In SPT034652, the polarised dust emission is offset from the total dust emission, similar to the polarised dust distribution in Antennae galaxies (Lopez-Rodriguez et al. 2022a), which is likely due to the disruption of ordered fields by the merger-trigger starbursts. In addition, the bimodal distribution of also favour the merger scenario. The position angles of the field lines depend only on the ratio of Stokes and , which have received identical amplification and averaging from the strong lensing. The presence of two groups of in the image plane indicates that the source plane should also have at least two magnetic field components, i.e., and in Fig. 2. For a merging system like the Antennae galaxies, several configurations of magnetic field can be formed, i.e. the existing fields on each galaxy, the fields connecting both galaxies, and the newly stretched lines in the tidal tail. Alternatively, a disk could also show a bimodal distribution of , but it needs to be highly inclined towards the sightline (such as NGC7331 in Borlaff et al. 2023), and it cannot explain why the polarised dust mostly overlapped with a single velocity component. Besides, the dynamo predicts regular magnetic fields with axisymmetric position angle in the disk, which is not fully consistent with the field structures from the lens reconstruction, where ordered fields have only been observed in the north and east sides of the lensed galaxy. The compact size of SPT034652 favours a late-stage merger. It supports the possibility that galactic magnetic fields have already been established in the two merging galaxies in SPT034652, and the merger-triggered starburst makes the B-field more turbulent and leaves mainly the ordered magnetic fields in the outskirts, such as the north and east sides of the lensed galaxy in the source plane. Therefore, the galaxy merger provides the most natural explanation for the formation of ordered magnetic fields in SPT034652.
4.2.2 Outflow driven?
@@
A strong outflow is present in SPT034652 based on the high-resolution dust continuum and H2O() observations (Jones et al. 2019a). It has a velocity up to 500 km s-1 and a substantial outflow rate, 100-900 M⊙ yr-1, likely driven by the galactic wind powered by the compact starbursting core. The galactic wind can distribute amplified magnetic fields to a larger volume (Brandenburg et al. 1993; Arámburo-García et al. 2021), thus can also explain the ordered field in the outskirts of SPT034652.lactic disks (Lopez-Rodriguez et al. 2021).
@@
However, the outflow can only partially explain the observations. Firstly, outflow can provide a distinct component along with the components from the host galaxy (Lopez-Rodriguez 2023), which can easily explain the bimodal distribution of in SPT034652, but the recovered field lines in the source plane seem to disfavour the bipolar configuration if SPT034652 is a disk galaxy. Besides, while the outflow scenario is compatible with the ordered field overlapped with a single velocity component, the relatively low-velocity dispersion and velocity gradient in the region with polarization detection also disfavours the outflow as the main driver. Future higher resolution follow-up of SPT034652 could provide better constraints.
4.2.3 AGN driven?
@@
It is also worth noting the possible contribution from the AGN in SPT034652. Powerful jets from AGN play an important role in magnetising the intracluster medium (ICM, Dubois et al. 2009; Xu et al. 2009). Meanwhile, the large-scale shocks induced by the jet and the AGN-driven outflow could also produce large-scale shear to order the turbulent fields. In the less extreme radio-quiet AGNs, the ubiquitous presence of compact radio jets (e.g. Girdhar et al. 2022; Rivera et al. 2023) also indicates their possible contribution to the development of magnetic fields in the ISM. In SPT034652, however, the presence of AGN is still elusive. The spectrum energy distribution (SED) analysis suggests a negligible AGN contribution to the total energy output, which is supported by the non-detection of X-ray and faint radio emission (Ma et al. 2016). In nearby AGN host galaxies, only the radio-loud AGN show highly polarised FIR fraction (5–11 per cent) in the core region; the radio-quiet AGN are mostly unpolarised (Lopez-Rodriguez et al. 2023). Therefore, even though there is a possible embedded AGN in SPT034652, our results support that it plays a negligible role in ordering the galactic magnetic fields. Future surveys targeting the AGN host galaxies are needed to test AGN’s importance in magnetising the ISM.
@@ 4.2.4 Caveats
In addition to the physical origins, the spatial resolution could also affect the observed polarisation. For the FIR polarisation, the polarised dust emission is density weighted by each individual molecular cloud, whose size is well below the current observation. A coarser spatial resolution thus introduces stronger beam averaging, which could counterbalance the polarised signal if the field lines are highly tangled or change orientation rapidly within each beam (Girichidis 2021). Gravitational lensing further complicates the situation by giving a different physical resolution along and perpendicular to the critical curve. Nevertheless, it represents the best way to mitigate the beam averaging. A robust polarisation signal has been observed in 9io9 (Geach et al. 2023), where the observation also benefits from a large lens amplification () to probe the dust polarisation down to 600 pc. In SPT034652, the regions with the polarisation are also those with the highest amplification, which indicates the non-detection in the other regions could be attributed to a combination of more turbulent magnetic fields and stronger beam averaging. This is also likely to be the reason for the non-detection in two other unlensed submm galaxies observed by Kade et al. (2023), where the physical resolution is worse than kpc.
@@ 4.3 Formation timescales of galactic magnetic fields
Theoretically, the amplification via small-scale dynamos is efficient, capable of amplifying the seed magnetic fields to G turbulent magnetic within a timescale of 100 Myr (Schleicher et al. 2010; Pakmor et al. 2014; Rieder & Teyssier 2016; Martin-Alvarez et al. 2022). However, the large-scale dynamos may need timescales of several Gyr to create kpc-scale ordered fields (e.g. Wang & Abel 2009; Kotarba et al. 2009; Arshakian et al. 2009; Rodrigues et al. 2019).
@@
The starburst nature of SPT034652 allows efficient small-scale dynamos in the ISM. In local star-forming galaxies, the total magnetic field strength correlates with the SFR surface density, i.e. with to be around 0.3 (e.g. Heesen et al. 2014), indicating a close bond between the magnetic fields and star formation activities. In SPT034652, the SFR surface density is about . Such concentrated SFR normally indicates a highly turbulent ISM (Elmegreen & Scalo 2004; Bournaud et al. 2010), which may lead to the efficient amplification of turbulent magnetic fields. Meanwhile, the presence of turbulent magnetic fields has been used to connect the two SFR tracers: the total FIR luminosity and non-thermal radio luminosity (i.e. FIR-radio correlation) (Condon 1992; Yun et al. 2001; Bell 2003). The FIR-radio correlation has been observed up to by stacking analysis and, interestingly, it shows only mild variations with the redshift (e.g. Ivison et al. 2010; Magnelli et al. 2015; Delvecchio et al. 2021; Yoon 2024). This close correlation suggests that amplified magnetic fields may already be present in the star-forming galaxies up to , which supports the claim that small-scale dynamo efficiently amplifies the turbulent magnetic fields in the early star-forming galaxies. Fig. 4 shows the magnetic field strength as a function of redshift and total SFR. Due to the selection effects, the current two measurements of dust polarisation at , SPT0346-52 and 9io9, are both extreme starburst galaxies, which biased our results to the most extreme SFGs in the early Universe.
@@
As we have mentioned before, the excess of Faraday rotation in the quasar absorber and fast radio burst support the presence of regular magnetic fields in galaxies up to (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Mannings et al. 2023). The detection of dust polarisation in 9io9 shows the ordered magnetic fields are present at z=2.6 (Geach et al. 2023). The redshift of SPT0346-52 is 5.6, thus the detection of the dust polarisation indicates the development of kpc-scale ordered galactic magnetic fields can be less than 1 Gyr, which puts a strong constraint on the ordering timescale. The current results do not allow us to directly test the dynamo. However, if the stretching and compressing during a merger is the main reason for the observed ordered magnetic fields (see the discussion in section 4.2), it indicates the two merging galaxies may have already developed ordered fields before the final merger, which indicates the ordering process is already ongoing during the first Gyr of the cosmic history. Therefore, our results suggest that the ordering process from large-scale dynamos could be efficient in early galaxies. Similar to the Faraday rotation, a more robust sample covering different galaxy types, redshifts, and environments is required to confirm this claim.
@@
The detection of magnetic fields at Cosmic Dawn can also constrain the amplification of their strength. However, estimating the strength from dust polarisation is non-trivial. The most widely used method, the DavisChandrasekharFermi (DCF) method (Davis & Greenstein 1951; Chandrasekhar & Fermi 1953), has mostly been used in the Galactic molecular clouds and star-forming regions (Crutcher 2012; Pattle et al. 2022), where the spatial resolution is able to capture the turbulence of the small-scale magnetic fields. On the galactic scale, modified DCF has been applied to the outflow region of M82 and the central molecular zone of the Milky Way (Lopez-Rodriguez et al. 2021; Guerra et al. 2023), but all of them require the spatial resolution comparable to the turbulence scale. The physical resolution of our observation at the location with polarised dust emission has a physical resolution of pc in the source plane. However, the turbulent scale in starburst galaxies injected by supernovae is found to be about 50-500 pc (Elmegreen & Scalo 2004), which is smaller than our observation and thus prevents us from estimating the strength of magnetic fields reliably.
@@
As we detected dust emission polarisation, it indicates that there is an ordered magnetic field associated with the cold phase of the ISM. The maximum magnetic field strength that this cold phase could reach is the equipartition energy level with the turbulent kinetic energy of the cold phase. Therefore, we can derive an upper limit by assuming an equipartition between the turbulent kinetic and magnetic energies (see also Geach et al. 2023). The equipartition field strength is formulated as

### v3

@@ 4 Discussion
In Section 3, we reported the properties of polarised dust in SPT034652, which indicates that kiloparsec-scale ordered magnetic fields were already present at the Cosmic Dawn. Here, we focus on several key questions that can be addressed from these observations, including: (1) how different the magnetic fields are between the distant and local galaxies; (2) what the main drivers are in ordering galactic magnetic fields; (3) what the formation timescale is of galactic magnetic fields; (4) how large-scale early magnetic fields could affect galaxy evolution;
@@ 4.1 Polarisation fraction and morphology of magnetic fields
Owing to the wavelength coverage and spatial resolution provided by ALMA, we can observe similar rest-frame FIR wavelengths and reach comparable spatial resolutions to the dust polarisation observations in the nearby Universe, especially compared with what the High-resolution Airborne Wideband Camera Plus (HAWC+) has done in the SALSA Legacy Program (Lopez-Rodriguez et al. 2022b). In Fig. 4, we compare the measured dust polarisation fraction in SPT0346-52 with the measurements from the SALSA Legacy Program. The rest-frame wavelength probed in SPT034652 is at 131 m, averaged over a physical resolution of 700 pc, making it comparable to the spatial resolution of SALSA Legacy Program at 154 m (250 pc-1.5 kpc, Lopez-Rodriguez et al. 2022b). A typical mean resolved polarisation fraction of percent has been reported among all the galaxies in the SALSA survey, dropping to a lower fraction of percent when only starburst galaxies (SFR) are considered. We have also included the dust polarisation measurements of 9io9 (Geach et al. 2023). The similar polarisation fraction between SPT034652 and the local starbursts is surprising, even though there are differences of two orders of magnitude in SFR and of more than 12 Gyr in cosmic time.
In SPT034652, the polarised dust emission is more patchy. We only observe polarised dust emission in a part of the galaxy and, interestingly, it does not overlap with the peak of total dust intensity (see Fig. 2). In Fig. 5, we compare the spatial distribution of polarisation with the kinematics from [C ii]. The region with the polarisation detection is also the region with the lowest velocity gradient and velocity dispersion. In the channel maps of [C ii] emission, the polarisation is mostly associated with a single velocity component near . This result is different from the configuration of polarised dust emission in 9io9, where the polarised dust largely overlaps with the total dust emission and the inferred fields line are aligned with the cold molecular gas disk (Geach et al. 2023). Such a difference suggests a possible different ordering process (see more discussion in section 4.2).
@@ 4.2 Physical origin of the ordered magnetic field
Various processes could lead to the formation of large-scale ordered magnetic fields, including the dynamo in galactic disks (Beck et al. 1994; Beck 2007), the stretching and compressing of existing fields by galaxy mergers, and galactic inflows and outflows (Chyży & Beck 2004; Drzazga et al. 2011; Basu et al. 2017; Jones et al. 2019b; Lopez-Rodriguez et al. 2022a; Lopez-Rodriguez 2023; Ledos et al. 2024). In this section, we discuss several possible mechanisms to explain the observed magnetic fields in SPT034652.
4.2.1 Disk-driven versus merger-driven
@@
High-resolution spectral line observations of SPT034652 have resolved a large velocity gradient ( 600 km s-1) of the cold gas in the ISM (Spilker et al. 2015; Litke et al. 2019). Due to the compact size of SPT034652 in the source plane, it is still unclear whether it is a rotating disk or a major ongoing merger. Litke et al. (2019) show that a lensing reconstruction of the high-resolution [C ii] emission reveals two distinct components, both spatially and kinematically. However, based on the observation of water absorption, Jones et al. (2019a) conclude that a compact disk with a strong molecular outflow gives a better explanation. In this subsection, we discuss both scenarios in generating the ordered magnetic fields.
@@
The formation of disks is the most discussed channel to develop ordered galactic magnetic fields (Beck et al. 1994; Pakmor et al. 2017, 2024). The presence of a galactic disk can provide continuous differential rotation to support the dynamo (Beck et al. 2019). Initially, the dynamo was not favoured in the early Universe, as most of the early galaxies were thought to be irregular and turbulent. This picture may need to be revised, however, given the apparent prevalence of disks in the early Universe (Förster Schreiber et al. 2018; Rizzo et al. 2023; Kuhn et al. 2023, Lee et al. (in prep.)). If galaxies can develop disks efficiently during the first Gyr of cosmic time, it is reasonable to expect large-scale ordered magnetic fields in the disks of early galaxies (Arshakian et al. 2009; Geach et al. 2023).
@@
In SPT034652, the rich observations allow a quantitative diagnosis of the dynamo associated with a galactic disk. Following the simplified model by Arshakian et al. (2009), we can define the dynamo number:
@@
where is the scale height of the disk, is the angular velocity of the disk, and are the maximum velocity and velocity dispersion of the disk, respectively, and is the maximum radius of the disk. Within a galactic disk, the dynamo is effective only when the dynamo number, , is larger than the critical value, , which is close to 7 in Milky Way-like disk galaxies. The from our lens model is about 1.5 kpc (see also Fig. 3). Assuming a typical disk scale height of 300 pc (Geach et al. 2023), the is of the order of . Considering the complex velocity field (merger+outflows), it is not straightforward to estimate the without some biased assumptions. In the most extreme case, assuming the velocity gradient is solely produced by a galactic disk, the maximum rotational velocity will be km s-1. We adopted the median velocity dispersion of 140 km s-1 from the region with the lowest velocity gradient, which is also the region with polarisation detection (see also Fig. 5). Then, the is about 2.5, which predicts , well below the (Ruzmaikin et al. 1988; Arshakian et al. 2009, see also the discussion in Geach et al. (2023)). Therefore, even if there were a disk in SPT034652, it would not be very effective in ordering the galactic magnetic fields. However, if the dynamic cold disk () could exist at such an early cosmic time (e.g. Rizzo et al. 2020; Neeleman et al. 2020; Lelli et al. 2021; Rowland et al. 2024), the dynamo number could be large enough to allow for an efficient dynamo.
@@
Galaxy mergers can compress and stretch magnetic fields, creating large-scale ordered magnetic fields. This process has been well demonstrated in the local major merger prototype – NGC4038/39 (the Antennae) (Chyży & Beck 2004; Basu et al. 2017). In the cores of the two merging galaxies of the Antennae system, the field lines are mostly dominated by turbulent fields. However, in the tidal tail, the magnetic fields are regular up to kpc, which is likely generated by stretching the initial disk fields in the pre-merger galaxies (Basu et al. 2017). A 9 kpc ordered magnetic field is also reported by means of magnetically aligned dust grains, which again supports a compressed and stretched magnetic field due to the merger activity (Lopez-Rodriguez et al. 2022a). The presence of ordered magnetic fields has also been confirmed in a sample of interacting galaxies observed with synchrotron polarisation (at 4.86 and 1.4 GHz) (Drzazga et al. 2011), which all support tidal interactions being able of effectively magnetising the mergers’ surroundings and producing large-scale ordered field structures up to tens of kiloparsecs in the ISM of the merged galaxies. Besides, the developed kiloparsec-scale ordered magnetic fields have a strength several times larger than the magnetic fields in disks from non-interacting galaxies (Drzazga et al. 2011; Beck & Wielebinski 2013). Similar intense field amplification and creation of large-scale magnetic fields with galaxy mergers have also been noticed in magneto-hydrodynamic simulations, which is crucial for angular momentum transportation and subsequent morphological evolution of the merged galaxies (Whittingham et al. 2021, 2023).
@@
In SPT034652, the polarised dust emission is offset from the total dust emission, similar to the polarised dust distribution in Antennae galaxies (Lopez-Rodriguez et al. 2022a), which is likely due to the disruption of ordered fields by the merger-trigger starbursts. In addition, the bimodal distribution of also favours the merger scenario. The position angles of the field lines depend only on the ratio of Stokes and , which have received identical amplification and averaging from the strong lensing. The presence of two groups of in the image plane indicates that the source plane should also have at least two magnetic field components; that is and in Fig. 2. For a merging system like the Antennae galaxies, several configurations of magnetic field can be formed, i.e. the existing fields on each galaxy, the fields connecting both galaxies, and the newly stretched lines in the tidal tail. Alternatively, a disk could also show a bimodal distribution of , but it needs to be highly inclined towards the line of sight (such as NGC7331 in Borlaff et al. 2023), and it cannot explain why the polarised dust mostly overlaps with a single velocity component. Besides, the dynamo predicts regular magnetic fields with an axisymmetric position angle in the disk, which is not fully consistent with the field structures from the lens reconstruction, in which ordered fields have only been observed on the north and east sides of the lensed galaxy. The compact size of SPT034652 favours a late-stage merger. It supports the possibility that galactic magnetic fields have already been established in the two merging galaxies in SPT034652, and the merger-triggered starburst makes the B-field more turbulent and leaves mainly the ordered magnetic fields on the outskirts, such as the north and east sides of the lensed galaxy in the source plane. Therefore, the galaxy merger provides the most natural explanation for the formation of ordered magnetic fields in SPT034652.
4.2.2 Outflow-driven scenario
@@
A strong outflow is present in SPT034652 based on the high-resolution dust continuum and H2O() observations (Jones et al. 2019a). It has a velocity up to 500 km s-1 and a substantial outflow rate, 100-900 M⊙ yr-1, likely driven by the galactic wind powered by the compact starbursting core. The galactic wind can distribute amplified magnetic fields to a larger volume (Brandenburg et al. 1993; Arámburo-García et al. 2021), and thus can also explain the ordered field on the outskirts of SPT034652
@@
However, the outflow can only partially explain the observations. Firstly, the outflow can provide a distinct component along with the components from the host galaxy (Lopez-Rodriguez 2023), which can easily explain the bimodal distribution of in SPT034652, but the recovered field lines in the source plane seem to disfavour the bipolar configuration if SPT034652 is a disk galaxy. Besides, while the outflow scenario is compatible with the ordered field overlapped with a single velocity component, the relatively low velocity dispersion and velocity gradient in the region with polarization detection also disfavours the outflow as being the main driver. Future high-resolution follow-up of SPT034652 could provide better constraints.
4.2.3 The role of AGN
@@
It is also worth noting the possible contribution from the AGN in SPT034652. Powerful jets from the AGN play an important role in magnetising the intracluster medium (ICM, Dubois et al. 2009; Xu et al. 2009). Meanwhile, the large-scale shocks induced by the jet and the AGN-driven outflow could also produce large-scale shear to order the turbulent fields. In the less extreme radio-quiet AGNs, the ubiquitous presence of compact radio jets (e.g. Girdhar et al. 2022; Rivera et al. 2023) also indicates their possible contribution to the development of magnetic fields in the ISM. In SPT034652, however, the presence of an AGN is still elusive. The spectral energy distribution analysis suggests a negligible AGN contribution to the total energy output, which is supported by the non-detection of X-ray and faint radio emission (Ma et al. 2016). In nearby AGN-hosting galaxies, only the radio-loud AGNs show a highly polarised FIR fraction (5–11 percent) in the galactic core; while the radio-quiet AGNs are mostly unpolarised (Lopez-Rodriguez et al. 2023). Therefore, even though there is a possible embedded AGN in SPT034652, our results indicate that it plays a negligible role in ordering the galactic magnetic fields. Future surveys targeting AGN-hostting galaxies are needed to test the importance of AGNs in magnetising the ISM.
@@ 4.2.4 Caveats
In addition to the physical origins, the spatial resolution could also affect the observed polarisation. For the FIR polarisation, the polarised dust emission is density-weighted by each individual molecular cloud, whose size is well below the current observation. A coarser spatial resolution thus introduces stronger beam averaging, which could counterbalance the polarised signal if the field lines are highly tangled or change orientation rapidly within each beam (Girichidis 2021). Gravitational lensing further complicates the situation by giving a different physical resolution along and perpendicular to the critical curve. Nevertheless, it represents the best way to mitigate the beam averaging. A robust polarisation signal has been observed in 9io9 (Geach et al. 2023), where the observation also benefits from a large lens amplification () that probes the dust polarisation down to 600 pc. In SPT034652, the regions with the polarisation are also those with the highest amplification, which indicates that the non-detection in the other regions could be attributed to a combination of more turbulent magnetic fields and stronger beam averaging. This is also likely to be the reason for the non-detection in two other unlensed submillimetre galaxies observed by Kade et al. (2023), where the physical resolution is worse than kpc.
@@ 4.3 Formation timescales of galactic magnetic fields
Theoretically, amplification via small-scale dynamos is efficient and capable of amplifying the seed magnetic fields to G turbulent magnetic levels within a timescale of 100 Myr (Schleicher et al. 2010; Pakmor et al. 2014; Rieder & Teyssier 2016; Martin-Alvarez et al. 2022). However, the large-scale dynamos may need timescales of several Gyr to create kiloparsec-scale ordered fields (e.g. Wang & Abel 2009; Kotarba et al. 2009; Arshakian et al. 2009; Rodrigues et al. 2019).
@@
The starburst nature of SPT034652 allows for efficient small-scale dynamos in the ISM. In local SFGs, the total magnetic field strength correlates with the SFR surface density; that is, , with being around 0.3 (e.g. Heesen et al. 2014), indicating a close bond between the magnetic fields and star formation activities. In SPT034652, the SFR surface density is about . Such a concentrated SFR normally indicates a highly turbulent ISM (Elmegreen & Scalo 2004; Bournaud et al. 2010), which may lead to the efficient amplification of turbulent magnetic fields. Meanwhile, the presence of turbulent magnetic fields has been used to connect the two SFR tracers: the total FIR luminosity and non-thermal radio luminosity (i.e. FIR-radio correlation) (Condon 1992; Yun et al. 2001; Bell 2003). The FIR-radio correlation has been observed up to by a stacking analysis and, interestingly, it shows only mild variations with redshift (e.g. Ivison et al. 2010; Magnelli et al. 2015; Delvecchio et al. 2021; Yoon 2024). This close correlation suggests that amplified magnetic fields may already be present in the SFGs up to , which supports the claim that small-scale dynamos efficiently amplify the turbulent magnetic fields in the early SFGs. Fig. 4 shows the magnetic field strength as a function of redshift and the total SFR. Due to selection effects, the two galaxies for which we currently have measurements of dust polarisation at , SPT0346-52 and 9io9, are both extreme starburst galaxies, which biases our results to the most extreme SFGs in the early Universe.
@@
As we have mentioned before, the excess of Faraday rotation in the quasar absorber and FRB supports the presence of regular magnetic fields in galaxies up to (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Mannings et al. 2023). The detection of dust polarisation in 9io9 shows that the ordered magnetic fields are present at z=2.6 (Geach et al. 2023). The redshift of SPT0346-52 is 5.6; thus, the detection of dust polarisation indicates that the development of kiloparsec-scale ordered galactic magnetic fields can be less than 1 Gyr, which puts a strong constraint on the ordering timescale. The current results do not allow us to directly test the dynamo. However, if the stretching and compressing during a merger is the main reason for the observed ordered magnetic fields (see the discussion in Section 4.2), it indicates the two merging galaxies may have already developed ordered fields before the final merger, which indicates that the ordering process is already ongoing during approximately the first Gyr of cosmic history. Therefore, our results suggest that the ordering process from large-scale dynamos could be efficient in early galaxies. Considering the similar caveats with Faraday rotation, a more robust sample covering different galaxy types, redshifts, and environments is required to confirm this claim.
@@
The detection of magnetic fields at Cosmic Dawn can also constrain the amplification of their strength. However, estimating the strength from dust polarisation is non-trivial. The most widely used method, the DavisChandrasekharFermi (DCF) method (Davis & Greenstein 1951; Chandrasekhar & Fermi 1953), has mostly been used in the Galactic molecular clouds and star-forming regions (Crutcher 2012; Pattle et al. 2022), where the spatial resolution is able to capture the turbulence of the small-scale magnetic fields. On the galactic scale, a modified DCF method has been applied to the outflow region of M82 and the central molecular zone of the Milky Way (Lopez-Rodriguez et al. 2021; Guerra et al. 2023), but each of these requires the spatial resolution to be comparable to the turbulence scale. The physical resolution of our observation at the location with polarised dust emission is pc. However, the turbulent scale in starburst galaxies injected by supernovae is found to be about 50-500 pc (Elmegreen & Scalo 2004), which is smaller than our observation, preventing us from estimating the strength of magnetic fields reliably.
@@
As we detected dust emission polarisation, this indicates that there is an ordered magnetic field associated with the cold phase of the ISM. The maximum magnetic field strength that this cold phase could reach is the equipartition energy level with the turbulent kinetic energy of the cold phase. Therefore, we can derive an upper limit by assuming an equipartition between the turbulent kinetic and magnetic energies (see also Geach et al. 2023). The equipartition field strength is formulated as